OpenStudy (anonymous):

so im looking at this question's solutions but i cant figure out how one of the steps is done... going from: 9a^4/64 = 9/8 a to a^3 = 8 .... ???? what am i missing?

OpenStudy (anonymous):

\[9a ^{4}/64 = (9/8)a \rightarrow 9a ^{4}/64 = 9a/8 \rightarrow 9a ^{4} = 9a*64/8\rightarrow a ^{3}=8\]

OpenStudy (anonymous):

Just multiply both sides by \(\frac{8}{9 a}\).

OpenStudy (anonymous):

anwar, how do get the divided by sign in the equation toolbox instead of 8/9a?

OpenStudy (anonymous):

is there somewhere to just practice this kind of thing over and over?

OpenStudy (anonymous):

OR you can use the equation feature and type frac{}{}.

OpenStudy (anonymous):

oh okay, thanks! it is so much easier to use that sign when explaining stuff rather than / symbol or the div symbol in the toolbox

OpenStudy (anonymous):

Lol I agree!! :D